ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 157, 2007

Bell, J. P. and S. Gerhold On the positivity set of a linear recurrence sequence

333

Betancor, J. J., J. C. Fariña, T. Martínez and J. L. Torrea Riesz transform and g-function associated with Bessel operators and their appropriate Banach spaces

259

Depauw, J. Degree two ergodic theorem for divergence-free stationary random fields

283

Einsiedler, M. and T. Fisher Differentiable rigidity for hyperbolic toral actions

347

Fariña, J. C. See J. J. Betancor

Fisher, T. See M. Einsiedler

Gaitsgory, D. On de Jong’s conjecture  155

Gerhold, S. See J. P. Bell

Hertweck, M. and M. Soriano Parametrization of central Frattini extensions and isomorphisms of small group rings

63

Hinrichs, A. and H. H. Lee Duality of Fourier type with respect to locally compact abelian groups

209

Hirschorn, J. Random trees under CH  123

Hu, P.-C. and C.-C. Yang Subspace theorems for homogeneous polynomial forms

47

Kaimanovich, V., I. Kapovich and P. Schupp The subadditive ergodic theorem and generic stretching factors for free group automorphisms

1

Kapovich, I. See V. Kaimanovich

Klartag, B. and R. Vershynin Small ball probability and Dvoretzky’s theorem

193

Lee, H. H. See A. Hinrichs

Martínez T. See J. J. Betancor Nair, R. and M. Pollicott Pair correlations of sequences in higher

dimensions

219

Paajanen, P. M. On the degree of polynomial subgroup growth in class 2 nilpotent groups

323

Pollicott, M. See R. Nair

Schupp, P. See V. Kaimanovich Soriano, M. See M. Hertweck Suzuki, T. Browder’s type convergence theorems for one-parameter

semigroups of nonexpansive mappings in Banach spaces

239

Tibăr, M. The vanishing neighbourhood of non-isolated singularities 309

Torrea, J. L. See J. J. Betancor Vershynin, R. See B. Klartag Wootton, A. Defining equations for cyclic prime covers of the Rie-

mann sphere

103

Yang, C.-C. See Hu, P.-C.